# Conformalized Interactive Imitation Learning: Handling Expert Shift & Intermittent Feedback

### v1

@@ Appendix A Proof of Proposition 1
$q_{1}\in[0,B]$ by assumption, so the lemma is satisfied at $t=1$. Assume $q_{t}\in[-\alpha N_{t-1},B+(1-\alpha)N_{t-1}]$. Now, for $q_{t+1}$, we consider Case 1, where $\text{obs}_{t}=0$, which means that $q_{t+1}=q_{t}$, so $q_{t+1}$ lies within the range $[-\alpha N_{t-1},B+(1-\alpha)N_{t-1}]$. Since $N_{t}\geq N_{t-1}$, $q_{t+1}$ also lies within the larger range $[-\alpha N_{t},B+(1-\alpha)N_{t}]$, as desired. In Case 2 where $\text{obs}_{t}=1$, $q_{t+1}=q_{t}+\frac{\gamma_{t}}{p_{t}}(\text{err}_{t}-\alpha)$. If we represent $\eta_{t}:=\frac{\gamma_{t}}{p_{t}}$, we obtain the constant-feedback quantile tracking update: $q_{t+1}=q_{t}+\eta_{t}(\text{err}_{t}-\alpha)$, with variable $\eta_{t}$ instead of $\gamma_{t}$. Then, Lemma 1 in Angelopoulos et al. (2024b), with $N_{t}=\max_{1\leq r\leq t}\eta_{t}$ bounds $q_{t+1}$ within the range $[-\alpha N_{t},B+(1-\alpha)N_{t}]$. ∎
@@
A.1 Special Case: When $\gamma_{t}=p_{t}$
@@
Appendix B Intermmittent Adaptive Conformal Inference
@@
B.1 Intermittent Adaptive Conformal Prediction
@@
At $\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}\text{err}_{t}$ approaches $\alpha$, given $M_{\infty}=\text{min }\{p_{1},...,p_{\infty}\}\in(0,1]$. This guarantees IACI gives the $1-\alpha$ long-term empirical coverage frequency regardless of the underlying data generation process. The proof follows from Gibbs & Candes (2021) as is described in the next subsection (Section B.2).
B.2 Proof for Intermittent Adaptive Conformal Inference
@@
Appendix C Extended Experiments: Intermittent Quantile Tracking
We experiment with these four predictors because of their use in prior conformal literature (Angelopoulos et al., 2024a) and in order to ensure that under different base model conditions, IQT maintains coverage close to the desired level. (1) Autoregressive (AR) model with 3 lags, (2) Theta model with $\theta=2$ (Assimakopoulos & Nikolopoulos, 2000), (3) Prophet model Taylor & Letham (2018), and (4) Transformer model (Vaswani et al., 2017). Consistent with prior works, for all base models except for transformer, we retrain the base model after each timestep; for the transformer, we retrain every 100 timesteps. We set lookback window $k=100$ timesteps for the Google and Amazon stock price data, and $k=300$ for the Elec2 dataset.
@@ Appendix C Extended Experiments: Intermittent Quantile Tracking
C.1 Google Stock Price Dataset Results
@@
C.2 Elec2 Dataset Results
@@
Appendix D Further Examination: Simulated Experiments
@@
Figure 10: Under $p_{t}=0.5$ intermittent feedback, ConformalDAgger learns the shift and drift more quickly than EnsembleDAgger.
@@
Figure 11: Under $p_{t}=0.9$ intermittent feedback, both ConformalDAgger and EnsembleDAgger learn the shift and drift quickly.
Appendix E Implementation and Task Details: Real Experiments
E.1 Learning policy training details

### v2

@@ Appendix C Proof of Proposition 1
$q_{1}\in[0,B]$ by assumption, so the lemma is satisfied at $t=1$. Assume $q_{t}\in[-\alpha N_{t-1},B+(1-\alpha)N_{t-1}]$. Now, for $q_{t+1}$, we consider Case 1, where $\text{obs}_{t}=0$, which means that $q_{t+1}=q_{t}$, so $q_{t+1}$ lies within the range $[-\alpha N_{t-1},B+(1-\alpha)N_{t-1}]$. Since $N_{t}\geq N_{t-1}$, $q_{t+1}$ also lies within the larger range $[-\alpha N_{t},B+(1-\alpha)N_{t}]$, as desired. In Case 2 where $\text{obs}_{t}=1$, $q_{t+1}=q_{t}+\frac{\gamma_{t}}{p_{t}}(\text{err}_{t}-\alpha)$. If we represent $\eta_{t}:=\frac{\gamma_{t}}{p_{t}}$, we obtain the constant-feedback quantile tracking update: $q_{t+1}=q_{t}+\eta_{t}(\text{err}_{t}-\alpha)$, with variable $\eta_{t}$ instead of $\gamma_{t}$. Then, Lemma 1 in Angelopoulos et al. (2024), with $N_{t}=\max_{1\leq r\leq t}\eta_{t}$ bounds $q_{t+1}$ within the range $[-\alpha N_{t},B+(1-\alpha)N_{t}]$. ∎
@@
C.1 Special Case: When $\gamma_{t}=p_{t}$
@@
Appendix D Intermmittent Adaptive Conformal Inference
@@
D.1 Intermittent Adaptive Conformal Prediction
@@
At $\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}\text{err}_{t}$ approaches $\alpha$, given $M_{\infty}=\text{min }\{p_{1},...,p_{\infty}\}\in(0,1]$. This guarantees IACI gives the $1-\alpha$ long-term empirical coverage frequency regardless of the underlying data generation process. The proof follows from Gibbs & Candes (2021) as is described in the next subsection (Section D.2).
D.2 Proof for Intermittent Adaptive Conformal Inference
@@
Appendix E Extended Experiments: Intermittent Quantile Tracking
We experiment with these four predictors because of their use in prior conformal literature (Angelopoulos et al., 2023) and in order to ensure that under different base model conditions, IQT maintains coverage close to the desired level. (1) Autoregressive (AR) model with 3 lags, (2) Theta model with $\theta=2$ (Assimakopoulos & Nikolopoulos, 2000), (3) Prophet model Taylor & Letham (2018), and (4) Transformer model (Vaswani et al., 2017). Consistent with prior works, for all base models except for transformer, we retrain the base model after each timestep; for the transformer, we retrain every 100 timesteps. We set lookback window $k=100$ timesteps for the Google and Amazon stock price data, and $k=300$ for the Elec2 dataset.
@@ Appendix E Extended Experiments: Intermittent Quantile Tracking
E.1 Google Stock Price Dataset Results
@@
E.2 Elec2 Dataset Results
@@
Appendix F Further Examination: Simulated Experiments
On Expert Realizability. An outstanding challenge in the theory of DAgger is guaranteeing learning a high-quality policy from the expert when the expert is unrealizable. We assume, much like Spencer et al. (2021), that our expert is realizable. Whether or not this is true in practice with human experts is an open area of research. Nevertheless, similar to HG-DAggerKelly et al. (2019) and EnsembleDAggerMenda et al. (2019), our approach offers a practical approach to contend with these real-world challenges.
@@ Appendix F Further Examination: Simulated Experiments
Figure 11: Under $p_{t}=0.5$ intermittent feedback, ConformalDAgger learns the shift and drift more quickly than EnsembleDAgger and SafeDAgger, with fewer queries than LazyDAgger.
@@
Figure 12: Under $p_{t}=0.9$ intermittent feedback, all approaches learn the shift and drift quickly.
Appendix G Experiments in Robosuite Nut-Assembly Benchmark
We conducted additional experiments on Robosuite’s nut-assembly task, a benchmark for LazyDAgger Hoque et al. (2021b) and ThriftyDAgger Hoque et al. (2021a). For consistency, we used the LazyDAgger thresholds from Hoque et al. (2021a) and set the SafeDAgger threshold to match the safety value used in LazyDAgger. The nut-assembly task involved a hardcoded expert policy that sequentially rotates to pick up the nut, lifts it, moves it to the peg, and lowers it to place. We trained the base policy with 30 demonstration episodes.
We evaluated this task under two conditions: a No-Shift scenario, where the task remains unchanged during deployment, and an Environment-Shift scenario, where the peg location is altered during interactive deployment (highlighted in red in Figure 13).
Performance metrics included: (1) autonomous success rate—success rate in deployment without a human supervisor (50 rollouts), and (2) intervention-aided success rate—success rate with a human supervisor in the loop. During training, we also tracked the number of interventions, human actions, and robot actions per episode. These metrics were calculated only for successful episodes to avoid bias from maximum episode length, which can inflate action counts for less successful policies that frequently reach the time limit.
The task required grasping a ring in a random initial pose and threading it onto a cylinder at a fixed target location. This involved two key challenges motivating learning from interventions: (1) successfully grasping the ring, and (2) accurately placing it over the cylinder (Figure 13). A simulated human provided interventions by teleoperating the robot. The state inputs included the robot’s joint angles and the ring’s pose, while the actions involved 3D translation, 3D rotation, and gripper control. We used 30 publicly available offline demonstrations (https://github.com/ryanhoque/thriftydagger 2,687 state-action pairs) from a human supervisor to initialize the robot policy across all algorithms.
ConformalDAgger identified the initial policy’s nonperformance in both contexts, prompting more queries and achieving a higher autonomous rollout success rate compared to other baselines.
Figure 13: ConformalDAgger recognizes the learned policy from 30 demos is still nonperformant in both contexts, and increases the number of questions asked, resulting in a higher autonomous rollout success rate than other baselines.
Appendix H Implementation and Task Details: Real Experiments
H.1 Learning policy training details